OpenStudy (anonymous):

1/4x^3=2 (x-2)^3=-64 1/8x^6-3=5

OpenStudy (anonymous):

1) \[\frac14 x ^{3} =2\rightarrow x ^{3} = 2*4\rightarrow x ^{3} =8\] \[ x = \sqrt[3]{8} \rightarrow x =\]

OpenStudy (anonymous):

oops sorry x = 2

OpenStudy (anonymous):

this is too frustrating to solve on OpenStudy's equation editor, so I'm going to solve it in GoogleDocs and then attach it. Be right back.